how to solve x2y''-3xy'-2y=0


1
Expert's answer
2021-06-28T16:37:01-0400
"x^2y''-3xy'-2y=0"

"y=x^m , y'=mx^{m-1}, y''=m(m-1)x^{m-2}"


"x^2(m(m-1)x^{m-2})-3x(mx^{m-1})-2x^m=0"


"(m^2-m-3m-2)x^m=0"


"(m^2-4m-2)x^m=0"

"m^2-4m-2=0"

"m=2\pm\sqrt{6}"

"y(x)=C_1x^{2-\sqrt{6}}+C_2x^{2+\sqrt{6}}"